OpenStudy (anonymous):

Please help! Step by step on how to solve sqrt(6t+3)^5

OpenStudy (anonymous):

\[\sqrt{(6t+3)^5}\]

OpenStudy (agent0smith):

Maybe this will help \[\Large \sqrt{(6t+3)^5} = \sqrt{(6t+3)^2*(6t+3)^2*(6t+3)}\]

OpenStudy (anonymous):

there is nothing to solve. Unless you meant simplify. But then, there is nothing to simplify either

OpenStudy (anonymous):

[\Large \sqrt{(6t+3)^5} = \sqrt{(6t+3)^2*(6t+3)^2*(6t+3)}\] First you do this then ,then you multiply 6t+3 then last you solve it.

OpenStudy (agent0smith):

It can be simplified @sourwing although not very much, \[\Large \sqrt{(6t+3)^5} = \sqrt{(6t+3)^2*(6t+3)^2*(6t+3)} = \] \[\Large (6t+3)^2\sqrt{(6t+3)} \]

OpenStudy (anonymous):

@agent0smith this is more like an example for those who learn roots to practice doing root. But in my view, (6t+3)^2 sqrt(6t+3) doesn't look any simpler than [sqrt(6t+3)]^5.

OpenStudy (anonymous):

Thanks guys so much! I appreciate the help and discussion

OpenStudy (agent0smith):

It doesn't look simpler @sourwing but I'm pretty sure that mathematically, it is simplified (assuming the question asked to simplify)

OpenStudy (anonymous):

@IloveCharlie but take @agent0smith 's answer. I'm sure that is what the question is asking you to do

OpenStudy (agent0smith):

Also, in the "simplified" form, you could add other terms with a common square root (of 6t+3)